HUBBLE SPACE TELESCOPE DAILY REPORT #843 PERIOD COVERED: 0800 EST Daily Status Report as of 074/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Three Sets of FOS/RD 4698 (FOS Dark Monitoring) The Faint Object Spectrograph was used to take a series of three internal flat exposures to determine the instruments background. All three observations were reported to look as expected. 1.2 Completed Two Sets of HSP/POL 3557 (Polarimetric Observations of the Crab Pulsar in the UV) The High Speed Photometer was used to make a total of ten observations in two separate sets of the Crab Nebula Pulsar. The first three observations were received and displayed. They were reported to look okay. The fourth observations was lost because the downlink was rejected. The next three observations were received in real time and looked okay, however, there was one FGS recentering event. In the last three observations, one was reported to be okay, one had a FGS recentering event, and the last was lost because the downlink was rejected and no data was received. 1.3 Completed Two Sets of WFC 1306 (GTO Low Latitude Parallel Program) The Wide Field Camera was used to make Three low latitude images which were taken in parallel with the prime science observations. The first two observations were reported to look good. The third observation was reported to have a very bright satellite trail crossing the field of view. 1.4 Completed PC 4347 (WF/PC Calibration: Flood Retention Test) The Planetary Camera was used to take three Earth flat observations using filter F439W. The observations occurred during LOS, and still remain of the science tape recorder. They will be analyzed later. No problems were reported. 1.5 Completed WFC/PC 4109 (WF/PC Calibration: Rapid Internal Monitor) Both the Planetary Camera and the Wide Field Camera were used to make internal flat observations using filter F555W. This was to check for new forms of contamination (i.e., measles and daisies), and the return of quantum efficiency hysterises. The observation were executed as scheduled , and no problems were reported. 1.6 Uplinked Updated Ephemeris Tables, Per MOSES 2871 1.7 Completed FOC 4800 (Stellar Content of Wolf-Rayet Galaxies) The Faint object Camera was used to make one ultraviolet observation of the Wolf-Rayet galaxy IIZW40. The exposure completed without incident under course track using a single guide star. The images were reported to show only two point sources. 1.8 Completed GHRS/2 2536 (Deuterium in the Local Interstellar Gas) The Echelle High Resolution Spectrograph was used to make observations of interstellar absorption lines by viewing early type B class star HD52089. The baseline guide star acquisition failed. The GHRS was idle with the take data flag being down and no image was produced. An attitude reference update was uplinked from the ground, and the following guide star reacquisition got course track. The following observations were reported to show a nice spectrum. However, it is not known if the correct target was really centered in the aperture. The last two observations did not occur, because the previous exposure did not finish in time. 1.9 Completed FOC/96 4804 (UV Imaging of Nearby Galaxies with FOC/96) The Faint Object Camera (f/96) was used to make one ten minute snapshot using filter F220W. About 60% of the image was received in real time before the spacecraft went into LOS. The part of the image which was received was reported to look okay. 1.10 Completed FOS/RD 2563 (Supernova Intensive Study) The Faint Object Spectrograph was used to make spectral observations to the supernova in the Large Magellanic Cloud (aka, LMC-SN-1987). The guide star acquisition and the subsequent target acquisition failed. The following observations did show some weak features. A later set of observations were received and displayed. They were reported to show nice emission features in spite of the guide star acquisition problems. 1.11 Initiated FOS/RD 4112 (Evolution of Alpha Lyman and C IV Absorption Systems) The Faint Object Spectrograph was used to initiate observations of Quasar PG1216+069. The observations are currently under way, and will be analyzed later. 1.12 Initiated PC 4714 (WF/PC Calibration: PC Internal Delta Flats) The Planetary Camera was used to initiate internal flat observations which will be used for the restoration of Earth flat observations. The observations are currently under way, and will be analyzed later. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 8 Successful : 6 Acq errors due to large attitude errors. (CT 073/0834Z, and CT at 073/1651Z) Scheduled Reacquisitions : 25 Successful : 25 Losses of Lock : 0 2.2 FHST Updates: Scheduled : 19 Successful : 19 2.3 Significant HSTARs: None 2.4 Operations Note: None 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete FOS/RD 4112 (Evolution of Alpha Lyman and C IV Absorption Systems) 3.2 Complete PC 4714 (WF/PC Calibration: PC Internal Delta Flats) 3.3 Complete FOC/96 4804 (UV Imaging of Nearby Galaxies with FOC/96) 3.4 Complete PC 4714 (WF/PC Calibration: PC Internal Delta Flats) 3.5 Complete Two Sets of GHRS/2 3496 (Ultraviolet Atlas of Sirius) 3.6 Complete FOS 4698 (FOS Dark Monitoring) 3.7 Complete PC 4249 (Black Holes, Stellar Dynamics and Populations) 3.8 Complete PC 4712 (WF/PC Calibration Measles Monitor) 3.9 Complete GHRS/2 2537 (Physical Conditions of Interstellar Gas) 3.10 Initiate FOS/BL 3791 (Quasar Absorption Line Survey) 3.11 Initiate WFC 4713 (WF/PC Calibration: WFC Internal Delta Flats) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. -END-